# TeleRAG: Efficient Retrieval-Augmented Generation Inference with Lookahead Retrieval

### v1

@@
Abstract
Retrieval-augmented generation (RAG) extends large language models (LLMs) with external data sources to enhance factual correctness and domain coverage. Modern RAG pipelines rely on large datastores, leading to system challenges in latency-sensitive deployments, especially when limited GPU memory is available. To address these challenges, we propose TeleRAG, an efficient inference system that reduces RAG latency with minimal GPU memory requirements. The core innovation of TeleRAG is lookahead retrieval, a prefetching mechanism that anticipates required data and transfers it from CPU to GPU in parallel with LLM generation. By leveraging the modularity of RAG pipelines, the inverted file index (IVF) search algorithm and similarities between queries, TeleRAG optimally overlaps data movement and computation. Experimental results show that TeleRAG reduces end-to-end RAG inference latency by up to 1.72× on average compared to state-of-the-art systems, enabling faster, more memory-efficient deployments of advanced RAG applications.
Equal contribution. Work done during internship at UW. Correspondence to: Chien-Yu Lin ¡cyulin@cs.washington.edu¿.
1 Introduction
@@
(b) Different scenarios for RAG and the illustration of the proposed lookahead retrieval mechanism. It prefetches relevant data for retrieval from CPU to GPU, overlaps data transfer with the pre-retrieval stage, and accelerates retrieval with GPU-CPU cooperation.
@@
Figure 1: (a) Illustrations of RAG pipeline stages. (b) Overview of TeleRAG and comparison to baseline systems.
@@
Retrieval-augmented generation (RAG) has emerged as a powerful technique to enhance large language models (LLMs) by integrating them with external databases[29, 13, 57]. During inference, RAG retrieves relevant content from external data sources, usually indexed as vector datastores, to mitigate issues such as hallucinations [63, 74, 50] and incorporate up-to-date or private information [36, 64].
@@
Modern RAG applications share two key characteristics. (1) RAG applications are built as modular pipelines as shown in Figure 1(a), where a single query undergoes multiple rounds of LLM calls and retrievals, each of which serves different functions to improve overall output quality. For example, query transformation rounds [27, 100, 71, 10] generally occur before retrieval to refine the user’s query with LLMs (pre-retrieval generation). (2) RAG’s datastores are typically large, supported by recent works demonstrating that increasing the size of the datastore positively affects the performance of RAG applications [64, 50, 15, 33, 77].
@@
These characteristics create significant challenges for efficient RAG inference, especially in latency-sensitive applications such as customer chatbots [18, 7, 83], financial analysis [62, 65], and emergency medical diagnosis [53, 31]. First, the latency of RAG systems is inherently increased due to multiple rounds of LLM generation and retrieval. Second, although GPU can speed up both LLM and retrieval, the combined memory demand of these processes often exceeds available GPU resources, making it expensive or infeasible to fully host both operations on GPU. Consequently, in local or small-scale deployments, which are common for RAG applications handling private or sensitive data, the retrieval datastore is frequently off-loaded to CPU memory to alleviate GPU constraints. However, while CPU off-loading resolves memory limitations, it significantly increases retrieval latency, diminishing overall system efficiency.
@@
Addressing the latency challenge, several recent works have been proposed to accelerate RAG inference. These approaches include FPGA-based retrieval acceleration [41], reuse of KV-cache [44, 60, 92], and speculative document retrieval [87, 98]. However, none of these works directly address the substantial memory demands associated with retrieval, which remains a significant bottleneck for efficient RAG deployments.
In this paper, we identify a key opportunity in the widely adopted inverted file index (IVF) [80] retrieval method. IVF reduces retrieval latency by pre-clustering the datastore into sub-clusters, limiting the search space to relevant clusters of the retrieval query during runtime. This approach also offers a way to reduces GPU memory usage, as only a subset of clusters are needed for each retrieval query. For example, after identifying relevant clusters, their data can be dynamically transferred from CPU to GPU memory and accelerate the similarity searches on GPU. However, this approach incurs high CPU-GPU data transfer overhead, keeping the overall retrieval latency still high [49] (see §3).
@@
Our proposal. To tackle latency and memory bottlenecks in RAG inference, we introduce a novel mechanism called lookahead retrieval. Lookahead retrieval predicts which subsets of IVF clusters will likely be accessed during the retrieval phase and proactively transfers them from CPU to GPU during the pre-retrieval generation stage.
The idea behind lookahead retrieval is based on the observation that queries before and after the pre-retrieval generation stage share substantial semantic overlap. Consequently, IVF clusters relevant to the initial input query are also likely to be relevant to the refined query. To validate this intuition, we rigorously analyze IVF cluster assignments across original queries and refined queries produced by the pre-retrieval generation step, evaluating on six representative RAG pipelines and three datasets. Our analysis confirms a strong similarity between these two query forms (see §3).
@@
Leveraging this insight, we propose TeleRAG, an efficient inference system designed to optimize RAG latency while minimizing GPU memory consumption. TeleRAG employs lookahead retrieval to preemptively load relevant IVF clusters onto the GPU, effectively hiding CPU-GPU data transfer overhead during concurrent LLM generation. As illustrated in Figure 1(b), this approach significantly reduces retrieval latency without exceeding GPU memory constraints, enabling efficient execution of RAG applications in resource-limited environments.
Although TeleRAG significantly reduces latency by processing prefetched IVF clusters on the GPU, some relevant clusters might still be missed during prefetching. To maintain retrieval accuracy, TeleRAG concurrently executes similarity searches for these missed clusters on the CPU. The results from CPU searches are seamlessly merged with GPU results, ensuring complete retrieval precision. Thanks to the high cluster similarity between the initial and refined queries, the number of missed clusters is typically small, allowing TeleRAG to achieve low retrieval latency through efficient CPU-GPU coordination.
@@
A key challenge in deploying lookahead retrieval is determining the optimal number of IVF clusters to prefetch. Prefetching too many clusters increases data-transfer overhead, whereas fetching too few clusters could result in higher retrieval latency due to increased CPU processing. To address this, we propose a profile-guided approach coupled with an analytical model that dynamically determines the ideal prefetch amount based on pipeline characteristics and hardware configurations. Our adaptive method consistently outperforms static configurations, delivering substantial latency improvements across diverse RAG setups and hardware environments.
Results summary. We evaluate TeleRAG using a 61GB Wikipedia-based datastore [5] across six popular RAG pipelines built with the Llama model family [82] (3B, 8B, and 13B). Remarkably, TeleRAG supports retrieval from a 61GB datastore alongside a Llama3-8B (16GB) LLM on a single RTX 4090 GPU (24GB memory), significantly outperforming the CPU retrieval baseline. Experiments conducted on an RTX 4090 and an H100 GPU [68] demonstrate speedups of up to 2.68x and 2.49x, with average speedups of 1.70x and 1.75x, respectively. These results highlight the capability of TeleRAG to efficiently handle large-scale RAG inference tasks within tight GPU memory constraints, confirming its practical value for resource-limited deployments.
@@
2 Background
Figure 2: Overview of RAG.
2.1 RAG
RAG is a technique that enhances the capabilities of LLMs by integrating them with information retrieval to generate more accurate and relevant text [29, 13, 57]. The core idea behind RAG is to augment the LLM with relevant information retrieved from a large corpus of documents, improving the LLM’s ability to answer questions without hallucinations [63, 74, 50] and generate contents based on up-to-date or private information [36, 64].
RAG workflow. A basic RAG workflow involves several phases, including data store building, retrieval, and interactions with LLMs [24, 34, 23, 88, 84, 70, 28]. In order to build a data store, raw data in various formats is cleaned, converted to plain text, and divided into chunks. These chunks are then encoded into vectors using an embedding model and stored in a vector index, enabling efficient searching based on similarity. When a user provides a query, it is converted into a vector using the same embedding model, and the most similar chunks from the indexed database are retrieved. For a large-scale index, approximate algorithms such as the inverted file index (IVF) [80] (see §2.2 for detail) are commonly used to accelerate the search process. The retrieved chunks, along with the original user query, are combined and given as a prompt to the LLM. The LLM then generates a response that relies on the information provided in the retrieved documents. This approach enables dynamic and informative responses in both single- and multi-turn dialogues [57].
Modularity in modern RAG applications. However, naively integrating document retrieval into LLM generation can cause several issues. For example, the retrieval often struggles to find relevant content and may select irrelevant or mismatched chunks [27, 100, 61], and a single retrieval may not provide sufficient context, necessitating multiple retrieval rounds [29].
To solve these issues, most state-of-the-art RAG models adopt a modular approach that employs multiple rounds of LLM calls and retrievals for a single query  [29, 26, 34, 28, 23, 88]. Typically, they have the following types of steps (shown in Figure 2):
  • Pre-retrieval generation, used to assess whether retrieval is needed or to generate queries for retrieval. An example of a pre-retrieval technique is query transformation [61, 37, 27, 102, 93, 73, 72, 100, 58], which reformulates the original query to make it clearer and more suitable for the retrieval task.
@@ 2.1 RAG
  • Post-retrieval generation, generates response based on user query and retrieved documents. It can also perform additional process such as summarization [39, 51] or reranking[105, 81] on the retrieved documents.
@@
By proceeding through these functions, RAG applications can deliver more precise and contextually appropriate responses, significantly enhancing the capabilities of LLMs for various applications [45, 29].
2.2 Vector Index and Inverted File (IVF)
@@
[…]
where $D$ is the dimensionality of the vector determined by the embedding model, and $d$ denotes the distance function, which is typically the L2-norm or inner product [46].
To improve search efficiency, the inverted file index (IVF) algorithm is widely used for large-scale vector databases due to its simplicity and effectiveness. IVF partitions the data store into clusters and restricts searches to the most relevant clusters, reducing the computational cost. To obtain the cluster assignment of each stored vector, it performs clustering algorithm such as $k$-means [66] and partitions the database into $N_{c}$ clusters:
@@
which involves sorting. In this way, IVF reduces search space and accelerates the retrieval process. Here, the hyperparameter $L$ from the first step is also referred as nprobe [25]. When nprobe is larger, the IVF retrieval will be more accurate as it search more clusters. However, the retrieve latency is longer as more computation and data are needed.
@@
Since the search process is highly parallelizable among each cluster and each vector, this search algorithm can be highly accelerated by GPUs. Open-source libraries offer efficient GPU implementations [46, 75]. However, IVF does not reduce the memory requirement for the index since the data for all clusters must be stored. As a result, IVF still requires a substantial GPU memory footprint, which becomes a bottleneck when GPU memory capacity is the constraint of RAG systems. Moreover, recent work reported that increasing the size of the data store positively affects RAG application performance [64, 50, 15, 33, 77], exacerbating this issue.
2.3 Latency-Sensitive RAG Use Cases
In this paper, we focus on reducing inference latency for a single query, a critical objective in latency-sensitive RAG applications. These applications span a range of real-world scenarios, such as customer chatbots [18, 7, 83], financial analysis [62, 65], autonomous driving [17, 94], and emergency medical diagnosis [53, 31], where users expect near-instantaneous responses. In each of these contexts, maintaining low latency per query is paramount for providing timely and accurate results.
Batching multiple queries, a common strategy for improving throughput, is not considered in this work. It is because many RAG applications involve user-specific private data [86, 55, 21, 89, 30] and serve relatively small user bases, making batching infeasible. For large-scale RAG services [2, 1, 10], although batching is feasible, it increases the latency for each individual queries [103, 97, 79] and therefore, is not ideal for the latency sensitive applications.
A straightforward way to reduce inference latency is to scale up GPU resources, allowing both the LLM and retrieval datastore to reside in GPU memory. However, this approach is often cost-prohibitive and could lead to significant hardware underutilization, particularly in the single-query scenarios that this work targets. Therefore, in this paper, we aim to improve single-query RAG latency without significantly increasing GPU memory usage, ensuring that both cloud-based and local deployments can satisfy stringent performance requirements without incurring excessive costs.
3 Analyzing Latency of RAG Pipelines
@@
In this section, we analyze state-of-the-art RAG applications and identify their underlying systems challenges in achieving low latency. To conduct the analysis, we construct a 61GB vector index with the FAISS library [25], and set the IVF clusters to 4096 following common practice [11]. We use Llama-3-8B [6] as the LLM, and run our analysis on an H100 GPU with 80GB memory. §5.2 provides details about our experimental setup.
3.1 End-to-end Latency of RAG Pipelines

### v2

@@
Abstract.
[…]
RAG inference, Acceleration, Prefetching.
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: submissionid: 164
1. Introduction
@@
(b) Illustration of TeleRAG’s lookahead retrieval mechanism and comparison to different RAG systems. TeleRAG prefetches relevant retrieval data from CPU to GPU, overlaps data transfer with the pre-retrieval stage, and accelerates retrieval with GPU-CPU cooperation.
@@
Figure 1. Overview of RAG pipeline stages and comparison of TeleRAG to baseline systems.
@@
Retrieval-augmented generation (RAG) has emerged as a powerful technique to enhance large language models (LLMs) by integrating them with external databases (Gao et al., 2023b; Asai et al., 2024c; Lewis et al., 2020). During inference, RAG retrieves relevant content from external data sources, usually indexed as vector datastores, to mitigate issues such as hallucinations (Mallen et al., 2023; Ram et al., 2023; Khandelwal et al., 2019) and incorporate up-to-date or private information (Izacard et al., 2023; Min et al., 2023).
@@
Modern RAG applications share two key characteristics. (1) RAG applications are built as modular pipelines as shown in Figure 1(a), where a single query undergoes multiple rounds of LLM calls and retrievals, each of which serves different functions to improve overall output quality. For example, query transformation rounds (Gao et al., 2023a; Zheng et al., 2023; Peng et al., 2023; Asai et al., 2024a) generally occur before retrieval to refine the user’s query with LLMs (pre-retrieval generation). (2) RAG’s datastores are typically large, supported by recent works demonstrating that larger datastore consistently improves the performance of RAG applications (Min et al., 2023; Khandelwal et al., 2019; Borgeaud et al., 2022; Hardt and Sun, 2023; Shao et al., [n. d.]).
@@
These characteristics create several significant challenges for efficient RAG inference, especially in latency-sensitive applications such as customer chatbots (Chemmagate, 2024; Akkiraju et al., 2024; Tung, 2024), financial analysis (Malec, 2024; MyScale, 2024), and emergency medical diagnosis (Klang et al., 2024; GM and Dhar, 2024). First, the latency of RAG systems is inherently increased due to multiple rounds of LLM generation and retrieval. Second, although GPUs can speed up both LLM generation and retrieval, the combination of these processes poses significant memory requirements, often exceeding available GPU resources (see §3.1). Hence, it is expensive or even infeasible to fully host both operations on GPUs.
Consequently, in local or custom deployments, which are common for RAG applications with private or sensitive data, the retrieval datastore is often offloaded to CPU memory to alleviate GPU memory constraints. However, while CPU offloading resolves memory limitations, it significantly increases retrieval latency (see §3.1), diminishing overall system efficiency. This memory pressure is not limited to local deployments; even for serving scenarios in data centers, where many concurrent requests are processed together and GPUs are plentiful, managing large index sizes still presents a significant challenge for achieving high throughput with latency service level objectives (SLOs). Allocating additional GPU memory to the retrieval data store directly reduces the memory available for the KV cache of the LLM serving system, limiting the effective batch sizes (Kwon et al., 2023). In contrast, CPU retrieval would increase the per-request latency by a lot, highlighting the need for techniques that reduce GPU memory consumption while preserving low latency.
@@
To address the latency challenge, several recent works have been proposed to accelerate RAG inference. These approaches include hardware accelerators for retrieval (Jiang et al., 2023d; Quinn et al., 2025), reuse of KV-cache (Jin et al., 2024a; Lu et al., 2024; Yao et al., 2024), and speculative document retrieval (Wang et al., 2024b; Zhang et al., 2024b). However, these works do not address the substantial memory demands associated with retrieval. Therefore, designing a memory-efficient system for low-latency RAG deployments is crucial.
In this paper, we argue that a promising direction to reduce GPU memory usage while maintaining low latency lies in improving the design of the retrieval index. Specifically, the inverted file index (IVF) (Sivic and Zisserman, 2003) reduces retrieval latency by pre-clustering data and limiting runtime search to relevant clusters. It offers a way to leverage GPU-accelerated search while not increasing memory consumption by dynamically transferring only the necessary clusters from the CPU to the GPU. However, this data transfer overhead is a key bottleneck, and needs to be minimized (Khan et al., 2024) (see §3.2).
@@
Our proposal. To tackle latency and memory bottlenecks in RAG inference, we introduce a novel mechanism called lookahead retrieval, which predicts the subsets of IVF clusters that will likely be accessed during runtime and proactively transfers them from the CPU to the GPU before the retrieval stage.
The key observation behind lookahead retrieval is that there is substantial semantic overlap between queries before and after the pre-retrieval stage. Namely, IVF clusters used by the initial input query, which is available well before the retrieval stage, are also likely to be relevant to the actual retrieval query (see §3.3).
@@
Using this insight, we propose TeleRAG, an efficient inference system designed to optimize RAG latency while minimizing GPU memory consumption. TeleRAG employs lookahead retrieval to preemptively load relevant IVF clusters onto the GPU, effectively hiding CPU-GPU data transfer overhead during concurrent LLM generation. As illustrated in Figure 1(b), this approach significantly reduces retrieval latency without exceeding GPU memory constraints, enabling efficient execution of RAG applications. To maintain retrieval accuracy while adopting lookahead retrieval, TeleRAG simultaneously searches for clusters that were not prefetched (i.e., mispredicted) from the CPU and merges the results. TeleRAG also supports batch and multi-GPU inference.
@@
A key challenge in deploying lookahead retrieval is determining the optimal number of IVF clusters to prefetch: Prefetching too many clusters increases data-transfer overhead, whereas prefetching too few could result in higher retrieval latency due to increased CPU processing. To address this, we propose a profile-guided approach coupled with an analytical model that dynamically determines the ideal prefetch amount based on pipeline characteristics and hardware configurations.
[…]
@@
2. Background
Figure 2. Overview of RAG.
2.1. RAG
RAG is a technique that enhances the capabilities of LLMs by integrating them with information retrieval to generate more accurate and relevant text (Gao et al., 2023b; Asai et al., 2024c; Lewis et al., 2020). The core idea behind RAG is to augment the LLM with relevant information retrieved from a large corpus of documents, improving the LLM’s ability to answer questions without hallucinations (Mallen et al., 2023; Ram et al., 2023; Khandelwal et al., 2019) and generate contents based on up-to-date or private information (Izacard et al., 2023; Min et al., 2023).
RAG workflow. A basic RAG workflow involves several phases, including data store building, retrieval, and interactions with LLMs (Dixit, 2024; Ilin, 2023; Databricks, 2024; Wei, 2024; Vijeaswaran et al., 2024; Pathway, 2024; Gao, 2024). In order to build a data store, raw data in various formats is cleaned, converted to plain text, and divided into chunks. These chunks are then encoded into vectors using an embedding model and stored in a vector index, enabling efficient searching based on similarity. When a user provides a query, it is converted into a vector using the same embedding model, and the most similar chunks from the indexed database are retrieved. For a large-scale index, approximate algorithms such as the inverted file index (IVF) (Sivic and Zisserman, 2003) (see §2.2 for detail) are commonly used to accelerate the search process. The retrieved chunks, along with the original user query, are combined and given as a prompt to the LLM. The LLM then generates a response that relies on the information provided in the retrieved documents. This approach enables dynamic and informative responses in both single- and multi-turn dialogues (Lewis et al., 2020).
Modularity in modern RAG applications. However, naively integrating document retrieval into LLM generation can cause several issues. For example, the retrieval often struggles to find relevant content and may select irrelevant or mismatched chunks (Gao et al., 2023a; Zheng et al., 2023; Ma et al., 2023), and a single retrieval may not provide sufficient context, necessitating multiple retrieval rounds (Gao et al., 2023b).
To solve these issues, most state-of-the-art RAG models adopt a modular approach that employs multiple rounds of LLM calls and retrievals for a single query  (Gao et al., 2023b; Fan et al., 2024; Ilin, 2023; Gao, 2024; Databricks, 2024; Wei, 2024). Typically, they have the following types of steps (shown in Figure 2):
  • Pre-retrieval generation, used to assess whether retrieval is needed or to generate queries for retrieval. An example of a pre-retrieval technique is query transformation (Ma et al., 2023; Jagerman et al., 2023; Gao et al., 2023a; Zhou et al., 2022; Ye et al., 2023; Press et al., 2023; Peng et al., 2024; Zheng et al., 2023; Liu, 2022), which reformulates the original query to make it clearer and more suitable for the retrieval task.
@@ 2.1. RAG
  • Post-retrieval generation, generates response based on user query and retrieved documents. It can also perform additional process such as summarization (Jiang et al., 2023b; Kim et al., 2023b) or reranking (Zhuang et al., 2023; Stewart and Linsdell, 2024) on the retrieved documents.
@@
By proceeding through these functions, RAG applications can deliver more precise and contextually appropriate responses, significantly enhancing the capabilities of LLMs for various applications (Jin et al., 2024b; Gao et al., 2023b).
2.2. Vector Index and Inverted File (IVF)
@@
$\text{k-argmin}_{i=0:N}d(x,y_{i}),$
where $D$ is the dimensionality of the vector determined by the embedding model, and $d$ denotes the distance function, which is typically the L2-norm or inner product (Johnson et al., 2019).
To improve search efficiency, the inverted file index (IVF) algorithm is widely used for large-scale vector databases due to its simplicity and effectiveness. IVF partitions the data store into clusters and restricts searches to the most relevant clusters, reducing the computational cost. To obtain the cluster assignment of each stored vector, it performs clustering algorithm such as $k$-means (Norouzi and Fleet, 2013) and partitions the database into $N_{c}$ clusters:
@@
which involves sorting. In this way, IVF reduces search space and accelerates the retrieval process. Here, the hyperparameter $L$ from the first step is also referred as nprobe (Douze et al., 2024). When nprobe is larger, the IVF retrieval will be more accurate as it search more clusters. However, the retrieve latency is longer as more computation and data are needed.
@@
Since the search process is highly parallelizable among each cluster and each vector, this search algorithm can be highly accelerated by GPUs. Open-source libraries offer efficient GPU implementations (Johnson et al., 2019; Rapidsai, 2022). However, IVF does not reduce the memory requirement for the index since the data for all clusters must be stored. As a result, IVF still requires a substantial GPU memory footprint, which becomes a bottleneck when GPU memory capacity is the constraint of RAG systems. Moreover, recent work reported that increasing the size of the data store positively affects RAG application performance (Min et al., 2023; Khandelwal et al., 2019; Borgeaud et al., 2022; Hardt and Sun, 2023; Shao et al., [n. d.]), exacerbating this issue.
2.3. Use Cases and Motivation
In this paper, we aim to achieve low-latency RAG inference without driving up GPU memory consumption, which is relevant for both local and data center settings. This goal is critical for efficiently serving a wide range of real-world applications where users expect near-instantaneous responses, thus imposing strict latency SLOs. Examples include customer chatbots (Chemmagate, 2024; Akkiraju et al., 2024; Tung, 2024; AWS, [n. d.]), financial analysis (Malec, 2024; MyScale, 2024), autonomous driving (Cai et al., 2024; Yuan et al., 2024), and emergency medical diagnosis (Klang et al., 2024; GM and Dhar, 2024).
On the one hand, there is significant demand for RAG applications that involve user-specific private data in local settings (Wang and Chau, 2024; Lam et al., 2023; Chung et al., 2023; Wutschitz et al., 2023; Ghodratnama and Zakershahrak, 2023). In such scenarios, the available GPU resources are usually limited, and the system serves a single request at a time. Therefore, CPU offloading is necessary when using a large vector index, which introduces substantial latency as we will show in the next section.
On the other hand, for large-scale RAG services in data centers (per, [n. d.]; gen, [n. d.]; Asai et al., 2024a), the goal is to serve many concurrent user requests cost-effectively, necessitating high system throughput. However, maximizing LLM serving throughput is often bottlenecked by the GPU memory available for the KV cache, which dictates the maximum possible batch size (Kwon et al., 2023). If a large retrieval datastore is loaded onto GPUs, it consumes memory that could otherwise be used for a larger KV cache, thereby limiting throughput. Hence, falling back to slow CPU retrieval to save GPU memory is a common option (Jiang et al., 2025), but this is not desirable for latency-sensitive applications due to latency overhead.
In both cases, a straightforward way to reduce inference latency is scaling up GPU resources to allow both the LLM and datastore to reside in GPU memory, which is usually cost-prohibitive. Therefore, in this paper, we aim to design an efficient RAG inference system that can satisfy latency SLOs without requiring a significant amount of GPU memory.
3. Analyzing Latency of RAG Pipelines
@@
In this section, we analyze state-of-the-art RAG applications and identify their underlying systems challenges in achieving low latency. To conduct the analysis, we construct a 61 GB vector index with the Faiss library (Douze et al., 2024), and set the IVF clusters to 4096 following common practice (Asai et al., 2023). We use Llama-3-8B (AI@Meta, 2024) as the LLM, and run our analysis on an RTX 4090 (24 GB memory) and H100 (80 GB memory). §5.2 provides more details about our experimental setups.
3.1. End-to-end Latency of RAG Pipelines
(a) Batch = 1.
(b) Batch = 4.